# Connections between Combinations Without Specified Separations and Strongly Restricted Permutations, Compositions, and Bit Strings

### v3

@@ 4 Connections with compositions
Four classes of $\mathcal{Q}$ whose associated digraphs possess common nodes have been investigated (Theorems 5–8 in [2]). One of these classes (Theorem 8 in [2]) always has at least two inner cycles; the simplest member of this class is $\mathcal{Q}=\{1,4\}$ (see Fig. 3). Another class (Theorem 7 in [2]) has just one inner cycle but always has the metatiles $C_{[q+1]}C_{[l]}^{2}$ and $C_{[q+1]}S^{a-1}C_{[2l]}S$, where $a=\left\lvert\mathbb{N}_{q}\setminus\mathcal{Q}\right\rvert$, which are of the same length (see Fig. 5 in [2]). The metatile-generating digraph of one of the remaining classes has $a$ inner cycles (Theorem 5 in [2]). The $a=1$ case leads to the following result.
@@
From the digraph (Fig. 8) it can be seen that the metatiles are $S$ and $C_{[q+1]}C_{[p]}^{j}S$ for $j=0,1,2,\ldots$ which are of lengths 1 and $q+1+jp$, respectively. ∎
@@
The following theorem originates from Theorem 6 in [2] and as it is a rather general theorem, it employs some extra notation that we now describe. We let $\theta$ be the bit string representation of $\mathcal{Q}$ whereby the $j$-th bit from the right in $\theta$ is 1 if and only if $j\in\mathcal{Q}$. By $\lfloor\theta/2^{b}\rfloor$ we mean discarding the rightmost $b$ bits in $\theta$ and shifting the remaining bits to the right $b$ places. We use $\mid$ to denote the bitwise OR operation.
@@
Figure 9: Digraph for tiling a board with squares and combs corresponding to $\mathcal{Q}$ specified in Theorem 20. $\sigma_{i}$ is the bit string corresponding to filling the first $i-1$ empty cells of a comb with squares and discarding the leading 1s.
@@
The digraph (Fig. 9, [2]) generates the metatiles $S$, $C_{[q+1]}S^{a}$, $C_{[q+1]}S^{i-1}C_{[p_{i}]}$ for $i=1,\ldots,a-1$, and $C_{[q+1]}S^{a-1}C_{[q-r]}^{j}S$ for $j\in\mathbb{Z}^{+}$. Their respective lengths are 1, $q+1$, $q+1+p_{i}$, and $q+1+j(q-r)$. As $q-r=p_{a}>p_{a-1}$, the lengths are all distinct. ∎
The OEIS sequences to which the above theorem applies along with the corresponding $\mathcal{Q}$ are as follows: A224809 {2,4}, A375981 {1,4,5}, A375982 {2,3,5}, A375983 {2,4,5}.
@@ 5 Connections with bit strings
In the rest of this section we show that the elements of $\mathcal{Q}$ concerned with the bijection given in Theorem 21 are always a well-based sequence and hence the generating function for the number of equivalence classes and the recursion relation for the number of equivalence classes whose set representation is of size $k$ are straightforward to obtain. The sequence $q_{1},q_{2},\ldots,q_{m}$, where $q_{i}<q_{i+1}$ for all $i$, is said to be a well-based sequence if $q_{1}=1$ and for all 2-partitions of $q_{j}$, for $j=2,\ldots,q_{m}$, at least one of the two parts is a $q_{i}$ [11, 22]. Thus, for example, $1,3,4$ is not a well-based sequence since $4=2+2$ and 2 is not in the sequence. In the present context, we can regard $q_{i}$ as being the elements of $\mathcal{Q}$ with $q_{m}=q$, its largest element. In the proof of Theorem 22, we will use the following definition which is equivalent aside from also classing the empty set as being well based: the sequence is well based if $a=\left\lvert\mathbb{N}_{q}\setminus\mathcal{Q}\right\rvert$ is zero or if, for all $i,j=1,\ldots,a$ (where $i$ and $j$ can be equal), $p_{i}+p_{j}\notin\mathcal{Q}$, where the $p_{i}$ are the elements of $\mathbb{N}_{q}\setminus\mathcal{Q}$. E.g., the only well-based sequences of length 3 are the elements of the sets $\{1,2,3\}$, $\{1,2,4\}$, $\{1,2,5\}$, and $\{1,3,5\}$.
@@
If $\mathcal{Q}$ is constructed from subword $\omega$ as described in Theorem 21 and $q$ is the largest element of $\mathcal{Q}$, then the elements of $\mathcal{Q}$ are a well-based sequence with the additional more general property that $\Lambda\notin\mathcal{Q}$ where $\Lambda$ is any linear combination (with integer coefficients) of the $p_{i}$.
@@
Suppose $\omega$ is the length-$l$ bit string $b_{l}b_{l-1}\cdots b_{2}b_{1}$. Then $p_{i},p_{j}\notin\mathcal{Q}$ iff $b_{n}=b_{n-p_{i}}$ and $b_{n}=b_{n-p_{j}}$ for $n\leq l$ with $n\geq p_{i}+1$ and $n\geq p_{j}+1$, respectively. Hence we also have $b_{n}=b_{n-p_{i}-p_{j}}$ for $1+p_{i}+p_{j}\leq n\leq l$ which implies that $p_{i}+p_{j}\notin\mathcal{Q}$ and hence the elements of $\mathcal{Q}$ form a well-based sequence. In an analogous way, it is easily seen that we have $b_{n}=b_{n-\Lambda}$ where $\Lambda$ is any integer-coefficient linear combination of the $p_{i}$ such that $0<\Lambda<n$ and hence $\Lambda\notin\mathcal{Q}$. ∎
@@
Figure 10: Digraph for tiling a board with squares and $\mathcal{Q}$-combs when the $q_{i}\in\mathcal{Q}$ are a well-based sequence. Node label $\sigma_{i}$ (for $i=1,\ldots,a$) denotes the bit string corresponding to filling the first $i-1$ empty cells of the $\mathcal{Q}$-comb with squares and discarding the leading 1s [2].
@@
We first deal with the $q=0$ case. Then $\mathcal{Q}=\{\}$ and so $\bar{c}=0$ and $c=1$ which leaves $\tilde{g}(x,y)=1/(1-x-xy)$. In this case, each binary word is in an equivalence class of its own and so $E^{(\omega)}_{n,k}=\tbinom{n}{k}$ since this is the number of length-$n$ binary words that contain $k$ 1s. The generating function for $\tbinom{n}{k}$ is $1/(1-x-xy)$. For $q>0$, since $E^{(\omega)}_{n}=E^{(\omega)}_{n,0}=1$ for $10\leq n\leq q-$ 1, and, by Theorem 21, $E^{(\omega)}_{n,k}=S^{\mathcal{Q}}_{n-q,k}$ for $n\geq q$ and from Theorem 1 $S^{\mathcal{Q}}_{n-q,k}=B^{\mathcal{Q}}_{n,k}$ for $n\geq q$, overall we have $E^{(\omega)}_{n,k}=B^{\mathcal{Q}}_{n,k}$. Hence $\tilde{g}(x,y)=G(x,y)$, the generating function such that the coefficient of $x^{n}y^{k}$ is the number of tilings of an $n$ board that contain $k$ combs. The digraph for tiling a board with squares and $\mathcal{Q}$-combs when the elements of $\mathcal{Q}$ are a well-based sequence is shown in Fig. 10 [2]. The digraph has a common node ($\sigma_{1}$) and the inner cycles ($S^{i-1}C_{[p_{i}]}$ for $i=1,\ldots,a$) have lengths $p_{i}$. There is a single common circuit ($C_{[q+1]}S^{a}$), which is of length $q+1$, and one outer cycle ($S_{[1]}$), which is of length 1. The common circuit and all the inner cycles contain a single comb. From Theorem 5 we then have
@@
Using the result that $c+\bar{c}=1+\sum_{i=1}^{q}x^{i}y=1+(x-x^{q+1})y/(1-x)$ to simplify the denominator then gives (10). ∎
The following corollary, which is a generalization of a result obtained previously (Corollary 1 in [2]; Theorem 2 in [11]), is then easily obtained from (11) by applying Corollary 7.
@@ 6 Discussion
Theorem 6, which we derived here in order to obtain the generating function for $S^{\{1,5\}}_{n}$, will be used elsewhere to obtain generating functions for $S^{\mathcal{Q}}_{n,k}$ for various other classes of $\mathcal{Q}$. As $\left\lvert\mathbb{N}_{q}\setminus\mathcal{Q}\right\rvert$ increases, the digraphs become too complicated for the application of general results analogous to Theorem 6. In such instances, the repeated use of Lemmas 2, 3, and 4 will yield the generating function, as was demonstrated in §3.1 for the $\mathcal{Q}=\{1,6\}$ case, although this becomes laborious for larger digraphs. Another approach is to employ a transfer matrix technique. For a digraph with $v$ nodes, the basic method gives the following expression for the generating function [21]:
@@
Acknowledgements
The author thanks the editors and two anonymous referees for their time and the detailed and helpful comments. He also thanks Alex Sáiz for testing the program on his 64-bit machine.

### v4

@@ 4.2 Compositions with parts drawn from infinite sets
Four classes of $\mathcal{Q}$ whose associated digraphs possess common nodes have been investigated [2, Theorems 5–8]. One of these classes (Theorem 8 in All24) always has at least two inner cycles; the simplest member of this class is $\mathcal{Q}=\{1,4\}$ (see Fig. 3). Another class (Theorem 7 in All24) has just one inner cycle but always has the metatiles $C_{[q+1]}C_{[l]}^{2}$ and $C_{[q+1]}S^{a-1}C_{[2l]}S$, where $a=\left\lvert\mathbb{N}_{q}\setminus\mathcal{Q}\right\rvert$, which are of the same length (see Fig. 5 in All24). The metatile-generating digraph of one of the remaining classes has $a$ inner cycles (Theorem 5 in All24). The $a=1$ case leads to the following result.
@@
From the digraph (Fig. 8) it can be seen that the metatiles are $S$ and $C_{[q+1]}C_{[p]}^{j}S$ for $j=0,1,2,\ldots$. These are of lengths 1 and $q+1+jp$, respectively. ∎
@@
The following theorem originates from Theorem 6 in All24 and as it is a rather general theorem, it employs some extra notation that we now describe. We let $\theta$ be the bit string representation of $\mathcal{Q}$ whereby the $j$-th bit from the right in $\theta$ is 1 if and only if $j\in\mathcal{Q}$ (we use the same bit string representation of $\mathcal{Q}$ in the program given in the Appendix). By $\lfloor\theta/2^{b}\rfloor$ we mean discarding the rightmost $b$ bits in $\theta$ and shifting the remaining bits to the right $b$ places (we employ the same notation in Theorem 21). We use $\mid$ to denote the bitwise OR operation.
@@
Figure 9: Digraph for tiling a board with squares and combs corresponding to $\mathcal{Q}$ specified in Theorem 20 [2]. In this figure and the next, node label $\sigma_{i}$ denotes the bit string corresponding to filling the first $i-1$ empty cells of the $\mathcal{Q}$-comb with squares and discarding the leading 1s.
@@
The digraph (Fig. 9) generates the metatiles $S$, $C_{[q+1]}S^{a}$, $C_{[q+1]}S^{i-1}C_{[p_{i}]}$ for $i=1,\ldots,a-1$, and $C_{[q+1]}S^{a-1}C_{[q-r]}^{j}S$ for $j\in\mathbb{Z}^{+}$. Their respective lengths are 1, $q+1$, $q+1+p_{i}$, and $q+1+j(q-r)$. As $q-r=p_{a}>p_{a-1}$, the lengths are all distinct. ∎
The OEIS sequences to which Theorem 20 applies along with the corresponding $\mathcal{Q}$ are as follows: A224809 {2,4}, A375981 {1,4,5}, A375982 {2,3,5}, A375983 {2,4,5}.
@@ 5.1 Binary word equivalence classes
In the rest of this section we show that the elements of $\mathcal{Q}$ concerned with the bijection given in Theorem 21 are always a well-based sequence and hence the generating function for the number of equivalence classes and the recursion relation for the number of equivalence classes whose set representation is of size $k$ are straightforward to obtain. The sequence $q_{1},q_{2},\ldots,q_{m}$, where $q_{i}<q_{i+1}$ for all $i$, is said to be a well-based sequence if $q_{1}=1$ and for all 2-partitions of $q_{j}$, for $j=2,\ldots,q_{m}$, at least one of the two parts is a $q_{i}$ [12, 23]. Thus, for example, $1,3,4$ is not a well-based sequence since $4=2+2$ and 2 is not in the sequence. In the present context, we can regard $q_{i}$ as being the elements of $\mathcal{Q}$ with $q_{m}=q$, its largest element. In the proof of Theorem 22, we use the following definition, which is equivalent aside from also classing the empty set as being well based: the sequence is well based if $a=\left\lvert\mathbb{N}_{q}\setminus\mathcal{Q}\right\rvert$ is zero or if, for all $i,j=1,\ldots,a$ (where $i$ and $j$ can be equal), $p_{i}+p_{j}\notin\mathcal{Q}$, where the $p_{i}$ are the elements of $\mathbb{N}_{q}\setminus\mathcal{Q}$. E.g., the only well-based sequences of length 3 are the elements of the sets $\{1,2,3\}$, $\{1,2,4\}$, $\{1,2,5\}$, and $\{1,3,5\}$.
@@
If $\mathcal{Q}$ is constructed from subword $\omega$ as described in Theorem 21, then the elements of $\mathcal{Q}$ are a well-based sequence with the additional condition that $\Lambda\notin\mathcal{Q}$ where $\Lambda$ is any linear combination (with integer coefficients) of the $p_{i}$.
@@
Suppose $\omega$ is the length-$l$ bit string $b_{l}b_{l-1}\cdots b_{2}b_{1}$. Then $p_{i},p_{j}\notin\mathcal{Q}$ iff $b_{n}=b_{n-p_{i}}$ and $b_{n}=b_{n-p_{j}}$ for $n\leq l$ with $n\geq p_{i}+1$ and $n\geq p_{j}+1$, respectively. Hence we also have $b_{n}=b_{n-p_{i}-p_{j}}$ for $1+p_{i}+p_{j}\leq n\leq l$. This implies that $p_{i}+p_{j}\notin\mathcal{Q}$ and hence the elements of $\mathcal{Q}$ form a well-based sequence. In an analogous way, it is easily seen that we have $b_{n}=b_{n-\Lambda}$ where $\Lambda$ is any integer-coefficient linear combination of the $p_{i}$ such that $0<\Lambda<n$ and hence $\Lambda\notin\mathcal{Q}$. ∎
@@
Figure 10: Digraph for tiling a board with squares and $\mathcal{Q}$-combs when the $q_{i}\in\mathcal{Q}$ are a well-based sequence [2].
@@
We first deal with the $q=0$ case. Then $\mathcal{Q}=\{\}$ and so $\bar{c}=0$ and $c=1$. This leaves $\tilde{g}(x,y)=1/(1-x-xy)$. In this case, each binary word is in an equivalence class of its own and so $E^{(\omega)}_{n,k}=\tbinom{n}{k}$ since this is the number of length-$n$ binary words that contain $k$ 1s. The generating function for $\tbinom{n}{k}$ is $1/(1-x-xy)$. For $q>0$, since $E^{(\omega)}_{n}=E^{(\omega)}_{n,0}=1$ for $10\leq n\leq q-$ 1, and, by Theorem 21, $E^{(\omega)}_{n,k}=S^{\mathcal{Q}}_{n-q,k}$ for $n\geq q$ and from Theorem 1 $S^{\mathcal{Q}}_{n-q,k}=B^{\mathcal{Q}}_{n,k}$ for $n\geq q$, overall we have $E^{(\omega)}_{n,k}=B^{\mathcal{Q}}_{n,k}$. Hence $\tilde{g}(x,y)=G(x,y)$, the generating function such that the coefficient of $x^{n}y^{k}$ is the number of tilings of an $n$ board that contain $k$ combs. The digraph for tiling a board with squares and $\mathcal{Q}$-combs when the elements of $\mathcal{Q}$ are a well-based sequence is shown in Fig. 10 [2]. The digraph has a common node ($\sigma_{1}$) and the inner cycles ($S^{i-1}C_{[p_{i}]}$ for $i=1,\ldots,a$) have lengths $p_{i}$. There is a single common circuit ($C_{[q+1]}S^{a}$), which is of length $q+1$, and one outer cycle ($S_{[1]}$), which is of length 1. The common circuit and all the inner cycles contain a single comb. From Theorem 5 we then have
@@
Using the result that $c+\bar{c}=1+\sum_{i=1}^{q}x^{i}y=1+(x-x^{q+1})y/(1-x)$ to simplify the denominator gives (10). ∎
The following corollary, which is a generalization of a result obtained previously [12, Theorem 2] (and also Corollary 1 in All24), is then easily obtained from (11) by applying Corollary 7.
@@ 6 Discussion
Theorem 6, which we derived here in order to obtain the generating function for $S^{\{1,5\}}_{n}$, will be used elsewhere to obtain expressions for generating functions for $S^{\mathcal{Q}}_{n,k}$ in terms of the elements of $\mathcal{Q}$ for various other classes of $\mathcal{Q}$. As $\left\lvert\mathbb{N}_{q}\setminus\mathcal{Q}\right\rvert$ increases, the digraphs become too complicated for the application of general results analogous to Theorem 6. In such instances, the repeated use of Lemmas 2, 3, and 4 yields the generating function, as we demonstrate in §3.1 for the $\mathcal{Q}=\{1,6\}$ case, although this becomes laborious for larger digraphs. Then a more direct approach is to employ a transfer matrix technique. For a digraph with $v$ nodes, the basic method gives the following expression for the generating function [22]:
@@
Given the simple form of (13), one might ask why we have gone to the trouble of introducing a lot of terminology concerning digraphs and deriving Theorems 5 and 6. There are two reasons. The first is that the determinant of the transfer matrix of a general digraph with an unspecified number of nodes (such as those shown in Figs. 9 and 10 in this article or Figs. 5 and 6 in All24) is not trivial to simplify to give the generating function; it is much easier to read off the lengths of the cycles and circuits in the digraph and plug them into the formula for $G(x,y)$. Once we have a general expression for $G(x,y)$ for a class of $\mathcal{Q}$, there is no longer any need to draw any digraphs (and deal with transfer matrices) for an instance of that class. Second, having concepts such as inner and outer cycles and common nodes can lead to quick proofs, as is the case for Lemma 18 in the present work and the main theorem in a forthcoming article [6].
7 Acknowledgments
The author thanks the two anonymous referees for their time and the detailed and helpful comments. He also thanks Alex Sáiz for testing the program on his 64-bit machine.